Currently browsing tag

magnification

TEM Status

December 3, 2013. The TEM is operational and the digital camera is attached and working. I need to calibrate the magnification and camera length next Monday morning. I should be able to start limited training later next week. The vibration isolation equipment will be installed during the week of Dec …

Yet another FESEM update

February 4, 2013. The FESEM is still not working properly.  We have to limp by until next week when Hitachi is sending 2 engineers to make the repair.  The scope can be used but may give problems.  If you see a black screen or a distorted screen try to go to …

TEM

TEM update

August 21, 2012. The TEM electron beam problem appears to be repaired. I have been able to achieve a stable 200kV beam ~10 times now so I believe the problem is repaired. I will work to achieve 300kV over the next few weeks. There is one problem that remains to …